Search
Text Size:
Smaller Text Normal Text Larger Text

Facebook page YouTube channel

BlueSky profile YouTube channel

P71-506: The new populations of gamma-ray bursts revealed by SVOM and Einstein Probe

PI: Daniele Bjorn Malesani, Johan Peter Uldall Fynbo, Dong Xu



ALFOSC/StanCam + NOTCam ToO program with override status

ToO triggers have override status.

The OB generator will be used to request all observations.
The OBs will automatically appear in the OB queue and OB generated scripts will be automatically placed in the directory ~/scripts/71-506/.

For the observer:

Go to the directory
  cd ~/scripts/71-506
List the available scripts
  ls -lrt
and execute the OB generated scripts requested.
If you cannot find the corresponding observing script, please (re-)compile it from the Queue interface.

In the case of NOTCam, make sure that differential skyflats are obtained for the camera and filter setup that is requested.

Remember to fill the ToO or SoftToO tracking logs of this program.


P71 tracking

Program Tracking
Proposal 71-506
Semester P72: 20251001 - 20260401
Principal Investigator Daniele Bjorn Malesani, Johan Peter Uldall Fynbo, DongXu
Programme The new populations of gamma-ray bursts revealed by SVOM and Einstein Probe
Instrument AL+NC+ST
Time Allocation 35h hard ToO, rest as SoftToO
Total Nights 70h
Type RRM+ToO+SoftToO
Notes Override. Soft triggers during Nordic service or Technical nights
Total Time Spent 14:40 (hh:mm)
Proposal Date Time Type Targets Files Comments Status
71-506 2025-10-01 00:35 ToO GRB251001B Observed with good seeing Observed
71-506 2025-10-07 01:33 ToO GRB251007A ALIj070147 - ALIj070155, ALIj070156 - ALIj070169 Observed with clear sky and seeing varying between 0.7'' and 1.0'' Observed
71-506 2025-10-09 00:55 ToO GRB251001B STIj090008 - STIj090016 Observed with clear skies and seeing around 0.7". For skyflats, take those taken in the morning. Observed
71-506 2025-10-13 01:36 ToO GRB251013C, Feige110 ALIj130066 - ALIj130089, ALIj130130 - ALIj130137 Observed in clear skies and occasionally through thin clouds and approx. 0.6 arcsec. seeing. Spectroscopy block stopped early (after 2 successful exposures) as agreed with the PI's. Observed
71-506 2025-10-14 00:57 ToO GRB251013C ALIj140085 - ALIj140094 Clear skies, and the seeing was around 1.3". Observed
71-506 2025-10-15 00:00 ToO Could not be observed due to bad weather. Failed
71-506 2025-10-18 00:56 ToO GRB251013C ALIj180056 - ALIj180065 Observed through thin clouds, sub-arc seeing. Observed
71-506 2025-10-19 00:57 SoftToO GRB251013C ALIj190236 - ALIj190245 Observed through thin clouds and seeing of approx. 0.8 arcsec. Observed
71-506 2025-10-21 00:29 ToO GRB251017A ALIj210059 - ALIj210063 Target observed with windy conditions and seeing around 2-2.5". Observed
71-506 2025-10-22 00:28 ToO GRB251017A ALIj220056 - ALIj220060 Target observed with clear skies and good seeing. Observed
71-506 2025-10-26 01:41 ToO GRB251013C, GRB251025B ALIj260002 - ALIj260006, ALIj260023 - ALIj260036 Observed under clear sky/thin clouds and seeing around 0.9". Observed
71-506 2025-10-27 00:19 ToO GRB251025B ALIj270078 - ALIj270080 Observed under clear sky and with seeing 1.4". Observed
71-506 2025-10-29 00:30 ToO GRB251025B ALIj290044 - ALIj290048 Observed at 22:30 with low airmass. Great seeing conditions and grey, clear sky Observed
71-506 2025-10-29 00:58 SoftToO GRB251013C ALIj290049 - ALIj290058 Observed with similar conditions as GRB_b. Observed
71-506 2025-10-30 00:30 SoftToO GRB251025B ALIj300083 - ALIj300087 Observed with great seeing clear skies and stable conditions. Observed
71-506 2025-11-01 01:03 SoftToO GRB251017A, GRB251025B ALIk010073 - ALIk010077, ALIk010088 - ALIk010096 Thin clouds, and the seeing was around 1.8". Observed
71-506 2025-11-02 01:05 ToO GRB251103A ALIk020381 - ALIk020392 Thin clouds, and the seeing was around 1.0". "grb251103a_alfosc_ima_rz" was executed for only 20 minutes at the PI’s request to switch to the counterpart (files ALIk020381 - ALIk020383). Observed
71-506 2025-11-03 00:08 SoftToO none NCIk030714 - NCIk030718 Interrupted the OB due to seeing being above the limit of 1.3". Observations failed. Failed

 

The above information is collected from the ToO/Mon reports that have been created by the observer.

Should any of the observations have been carried out using the OB system, you will find additional information HERE

Back to top Last modified: October 02 2025